OpenStudy (anonymous):

separate 846 into 3 parts so that the second part is twice the first part and the third part is triple the second part. which of the following eqations could be used to solve the problem? A.) 846=x+2x+3x B.) 846=x+2x+6x C.) 846/3=x+2x+6x

OpenStudy (anonymous):

846 = x + 2x + 6x 846 = 9x 846/9 = x 94 = x (1st part) 2x = 94(2) = 188 (2nd part -- twice the first part) 6x = 94(6) = 564 (3rd part -- triple the 2nd part) ANSWER : 846 = x + 2x + 6x

OpenStudy (anonymous):

thnks :)

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

:)